OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL! find the value of x in the triangle. A- 2 degree's B- 29 degree's C- 63 degree's D- 299 degree's

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (triciaal):

the sum of the angles in a triange add up to 180 180 = x + 90 + 61

OpenStudy (triciaal):

typo triangle

OpenStudy (anonymous):

can u explain a little further?

OpenStudy (triciaal):

3 numbers add to 180 you have 2 of the numbers 180 = x + 61 + 90 solve for x the other number

OpenStudy (triciaal):

if 2 + 4 + another number = 7 how would you know what is missing? 7 - (2 + 4) so the other number =1 2 + 4 + 1 = 7 same process

OpenStudy (anonymous):

okkkk

OpenStudy (triciaal):

what's your answer?

OpenStudy (anonymous):

um b?

OpenStudy (triciaal):

you got it (show some confidence) b